I not sure how to write the equation for this one the one i did was wrong.
To get to his parents' house, John must travel at a speed of 60 mph on land and then use a motorboat that travels at a speed of 20 mph in still water. John goes by land to a dock and then travels on a river against a current of 4 mph. He reaches his parent's home in 4.5 hours. The return trip takes him 3.5 hours. How far do his parents live from his house?

Answer by scott8148(5891)

the land part of the trip is the same, coming or going, so concentrate on the water

distance equals rate times time __ it takes an hour longer to go upstream

let t="return time on water" __ (20-4)(t+1)=(20+4)(t) __ 16t+16=24t __ 16=8t __ t=.5

so 12 mi [24(.5)] on water and 180 mi [60(3.5-.5)] on land